# Controllable Patching for Compute-Adaptive Surrogate Modeling of Partial Differential Equations

### v1

@@ 4 Experimental Results
    We compare CKM and CSM to SOTA non-patch based PDE surrogates including TFNO [68], FFNO [69], SineNet [70], and Transolver [71]. These comparisons demonstrate that flexible patch-based models are not only viable but often outperform non-tokenized architectures in next-step and rollout prediction tasks.
@@
    We test whether a single compute-elastic model is preferable to training multiple fixed-patch models. We train three fixed-patch models (patch sizes 4, 8, 16) for 200 epochs each (App. B.3), then train CKM and CSM models using the combined compute budget.
    This setup reflects a practical question: given a fixed training budget, is it better to train one flexible model or several static ones? All evaluations are performed on models trained under this constraint. After training, we compare the performance of CKM and CSM models against individual models trained at static patch sizes. If the compute elastic models can match or improve over the fixed patch baselines, this enables the novel creation of a broad class of flexible PDE surrogate eliminating the need to train and maintain separate fixed-patch models.
@@
    We perform the experiments described above on two transformer-based processors, namely the axial ViT [22] and the vanilla ViT [9] to demonstrate the architecture agnostic–ness of the CKM/CSM strategies. [9].
@@
    We assess how these flexible modules improve stability, prediction accuracy and mitigate artifacts in extended rollouts (Sec.4.3).
We additionally provide a set of ablation studies in App G. The input spatio-temporal sequence consists of six time frames, with the loss function optimized for next-step prediction.
Figure 3: Next-step prediction test VRMSEs vs number of tokens at inference of a 100M parameter model trained on four 2D datasets (left to right):shear_flow, turbulent_radiative_layer_2D, active_matter and  rayleigh_benard. Lower VRMSE is better. Token count is a proxy for required compute. This plot highlights the compute-accuracy tradeoff expanded more in sec. 4.2
4.1 Comparison to non-patch based SOTA models
Although our focus is not on non-patch architectures, Table 1 provides a reference comparison to state-of-the-art CNN/Neural-Operator surrogates. It demonstrates that flexible models are not only viable, but highly competitive with SOTA PDE surrogate architectures that do not use patch-based tokenization (e.g., TFNO [68], FFNO [69], Transolver [71], SineNet [70]). In fact, our models often significantly outperform these baselines. This result reinforces the potential of adaptive patch-based transformers to serve as competitive surrogates for PDE modeling, without sacrificing accuracy or flexibility. We did this architectural comparison for two complex physical systems, shear_flow  and turbulent_radiative_layer_2D  from the Well [20]. In the results of the sections to follow, we will focus mainly on comparisons between our flexible models vs. static patch baselines.
Model TFNO FFNO SineNet Transolver CSM + ViT (Ours) CKM + ViT (Ours)
Number of parameters 20M 7M 35M 11M 7M 100M 7M 100M
shear_flow 0.89 0.11 0.17 $>1$ 0.0762 0.0538 0.096 0.0531
turbulent_radiative_layer_2D 0.81 0.485 0.65 $>1$ 0.458 0.299 0.477 0.321
Table 1: 10-step rollout VRMSE comparison across patch-free models and our patch-based surrogate variants. We report scores for both 7M and 100M parameter versions of CSM and CKM.
4.2 Prediction accuracy and inference time patch/stride tunability
Dataset # Tokens Axial ViT (50M) Vanilla ViT (100M) The Well Baselines
CKM CSM f.p.s. CKM CSM f.p.s.
shear_flow 2048 0.0070 0.0070 0.0085 0.0055 0.0055 0.0067 CNext-U-Net
512 0.0098 0.0104 0.0124 0.0086 0.0088 0.0096 0.1049
128 0.0143 0.0134 0.0173 0.0120 0.0110 0.0140
turbulent_radiative_layer_2D 3072 0.153 0.178 0.170 0.133 0.144 0.143 CNext-U-Net
768 0.165 0.178 0.186 0.153 0.178 0.170 0.2269
192 0.211 0.211 0.237 0.200 0.202 0.223
active_matter 4096 0.0240 0.0230 0.0275 0.0192 0.0172 0.0210 CNext-U-Net
1024 0.025 0.028 0.033 0.0220 0.0210 0.0240 0.0330
256 0.0369 0.0367 0.0430 0.034 0.029 0.031
rayleigh_benard 4096 0.0277 0.0290 0.0361 0.0250 0.0250 0.0310 CNext-U-Net
1024 0.0370 0.0398 0.0442 0.0330 0.0350 0.0360 0.2240
256 0.0538 0.0522 0.0581 0.0470 0.0460 0.0470
supernova_explosion 4096 0.288 0.305 0.288 0.266 0.287 0.272 U-Net
512 0.370 0.380 0.392 0.360 0.380 0.387 0.3063
64 0.419 0.421 0.434 0.413 0.417 0.426
turbulence_gravity_cooling 4096 0.124 0.138 0.150 0.103 0.121 0.118 CNext-U-Net
512 0.180 0.197 0.194 0.164 0.182 0.179 0.2096
64 0.216 0.215 0.214 0.202 0.200 0.198
Table 2: Test VRMSE across multiple 2D and 3D PDE datasets for next-step prediction using CKM, CSM, and fixed patch size (f.p.s.) models. CKM/CSM models are trained once and evaluated across multiple token counts (i.e., patch/stride configurations), whereas each fixed p.s. model is trained separately. Results are shown for two transformer backbones: Axial ViT (50M) [22] and Vanilla ViT (100M) [9]. We also report the best outcomes from the Well benchmark suite [20].
We choose the VRMSE metric (App. B.4) as a measure of accuracy. In Table 2 and Fig. 3, we report VRMSE across a range of 2D and 3D PDE datasets from the Well [20]. Each model is evaluated at multiple token counts, which correspond to varying patch sizes in CKM or stride sizes in CSM —highlighting the impact of inference-time compression on predictive accuracy. Importantly, the fixed patch size (p.s.) models are separate networks trained separately at each patch size, while CKM/CSM are trained only once using random patch/stride schedules and evaluated without retraining. We show results for two different transformer architectures, Axial ViT [22] and vanilla full attention ViT [9]. Importantly, the goal is not to compare the two base architectures against each other, but to show that for both architectures, higher token count implies greater accuracy, and that CKM/CSM can enable a token count tunability at inference against the corresponding fixed patch baselines without losing any accuracy.
Across all datasets and token configurations, CKM and CSM consistently match or outperform their fixed-patch counterparts—despite being single, compute-elastic models. This includes both axial and vanilla attention processors, demonstrating the architecture-agnostic nature of our encoder/decoder modifications. Across all datasets, we observe a consistent trend: increasing the number of tokens at inference—by using smaller patch sizes (CKM) or strides (CSM)—leads to improved VRMSE, but it is more expensive computationally. This behavior is shown in both Table 2 and Fig.3. For example, in the active_matter  dataset, with a spatial resolution of $256^{2}$, reducing the patch/stride size from 16 to 4 increases token count from 256 to 4096, triples inference time (0.21s/step → 0.63s/step), and increases compute from 5 to 170 GFLOPs—but also reduces error by over 30%. Similarly, for our 3D datasets with a spatial resolution $64^{3}$, decreasing the patch size from 4 to 16, or increasing the tokens from 64 to 4096 leads to an increase in the inference speed per step by approximately a factor of 8 (0.11s/step→ 0.8s/step); but also improves VRMSE by a factor of $\sim$2.
These results reinforce that token count serves as a natural proxy for compute-accuracy tradeoff in patch-based PDE surrogates. Importantly, CKM and CSM allow a single model to be flexibly deployed at different compute budgets, without retraining, unlike fixed-patch baselines. We additionally evaluate performance across spatial frequencies using the binned spectral normalized mean squared error (BSNMSE). Across datasets, CKM and CSM consistently match or outperform fixed patch models in all frequency bands, highlighting their ability to preserve both coarse and fine-scale features. Further definitions, and visualizations are provided in App.  C.
Finally, as also stated before, while our main goal is not to benchmark against non-patch-based surrogates, Table 2 still includes the best results from FNO [72], TFNO [68], U-Net, and CNext-U-Net [73] from the Well benchmark suite [20]. We re-trained these models with tuned learning rates and report their best accuracy. Despite this, our flexible models consistently outperform these baselines across most tasks and inference budgets, demonstrating that flexible patch-based surrogates are not only viable but can exceed the performance of traditional operator-based or convolutional surrogates, and outperform a range of SOTA architectures on complex datasets as discussed in sec. 4.1.
To further test the modularity of our flexible modules, we integrate it into the encoder of the Continuous Vision Transformer (CViT) [63], a recent hybrid architecture which features a patch-based transformer encoder coupled with a novel, grid- based continuous decoder that directly queries spatiotemporal locations. Since the decoder in CViT does not use patch-based tokenization, we apply CKM only to the encoder. Since CViT is designed for 2D problems, we restrict our evaluation to 2D datasets, ensuring a fair and focused test of patch flexibility without altering any other model components. Results across four 2D PDE datasets show that a flexified CViT consistently outperforms its fixed-patch CViT counterpart for both next–step and rollout results. This experiment underscores that CKM can be selectively applied even within hybrid ViT models, further reinforcing the architecture agnostic-ness and modularity of these strategies. Full results are presented in App.D.
This ability to dynamically scale computational resources at inference is, to our knowledge, new in the PDE emulator literature. While large language models leverage inference-time optimizations such as entropy-based techniques, PDE surrogate models typically lack such flexibility. Our compute-elastic models address this gap by enabling adaptive patches and strides, allowing dynamic resource allocation to balance accuracy and efficiency at inference.
Figure 4: Step 44 rollout for the $D_{xy}$ field of the active_matter dataset.
4.3 Improved rollouts with compute elastic models
A key advantage of compute-elastic models is their ability to enable novel inference-time rollout strategies that are inaccessible to fixed-architecture surrogates. We introduce one such strategy: an alternating-patch rollout, where CKM and CSM models dynamically cycle through patch or stride sizes of 4, 8, and 16 at each prediction step. To our knowledge, this is the first demonstration in the PDE surrogate modeling literature where inference-time patching is not only treated as a tunable knob, but actively used to improve long-horizon prediction fidelity.
This strategy significantly reduces a persistent issue in patch-based architectures: patch artifacts—grid-like distortions that commonly arise in autoregressive rollouts, especially for fixed patching. These artifacts, prevalent across ViT variants (vanilla, axial, Swin), stem not from training instability or data quality, but from the limitations of static tokenization (App. E for more extended evidence).
Figure 4 demonstrates this phenomenon visually in the active_matter  dataset. We see significant checkerboard artifacts at rollout step 44 for the static patch model, which are largely stabilized by our alternating rollout strategy. Quantitatively, Table 3 shows the 10-step rollout VRMSE for all datasets. Across both axial and vanilla transformer backbones, compute-elastic models consistently outperform fixed patch models—often by large margins—demonstrating greater long-horizon stability. In fact, comparing these 10-step rollout values again with the non-patch SOTA models in Table 1 for relevant datasets, we find that from a purely architectural comparison perspective, our models often achieve SOTA outcomes on these complex tasks. Fig. 5 additionally shows a full trajectory evolution of the shear_flow dataset comparing fixed patch vs CKM/CSM. Clearly, the fixed patch 16 model develops prominent artifacts already by step 40, while CKM/CSM (CKM especially in this case) maintain a much more stable long trajectory. For additional rollout visualizations and metrics, refer to the App. F. We emphasize the powerful insight here that these improvements emerge without any additional training, enabled purely through dynamic inference-time patch modulation.
Figure 5: 100 step full trajectory rollout of shear_flow . Top to bottom: ground truth, fixed patch 16 model, CKM+vanilla ViT and CSM+vanilla ViT.
Dataset Axial ViT (50M) Vanilla ViT (100M)
CKM CSM f.p.s. $\Delta$ CKM CSM f.p.s. $\Delta$
shear_flow 0.082 0.089 0.143 +42.7% 0.057 0.053 0.107 +50.5%
turbulent_radiative_layer_2D 0.570 0.475 0.571 +16.8% 0.409 0.373 0.446 +16.4%
active_matter 0.390 0.384 0.640 +40.0% 0.351 0.359 0.370 +5.1%
rayleigh_benard 0.2159 0.166 0.217 +23.5% 0.215 0.140 0.2273 +38.4%
supernova_explosion 1.270 1.310 1.905 +33.3% 1.14 1.20 1.75 +34.9%
turbulence_gravity_cooling 0.680 0.667 0.860 +22.4% 0.527 0.559 0.77 +31.6%

### v2

@@ C.3 Training configuration - Axial ViT and Vanilla ViT experiments
    Hardware and parallelism. We train with a global batch size of 16 on 8 NVIDIA A100 80GB GPUs using PyTorch FSDP. No gradient accumulation.
@@
    Optimizer and regularization. We use Adam with learning rate $10^{-4}$ and weight decay $10^{-4}$, and set drop path to 0.1. We tuned the learning rate over $\{10^{-5},10^{-4},10^{-3}\}$.
@@
    Training loss. All models are trained with Normalized Mean Squared Error (NMSE), averaged over spatial dimensions and fields.
@@
    Prediction target. We use delta prediction, i.e., the model predicts the change in the field from the previous timestep. This affects training only; validation and test metrics are computed on reconstructed fields.
  5. 5.
    Positional encoding. We use RoPE (Su et al., 2021).
  6. 6.
    Training details: Each fixed-patch baseline ($p\in\{4,8,16\}$) is trained for 20,000 optimizer steps (200 epochs with epoch size of 100). The compute-elastic CSM/CKM model is trained with a matched total number of optimizer updates equal to the sum of the three fixed-patch runs, sampling the stride/patch setting uniformly from $\{4,8,16\}$ at each update (same global batch size and hardware). Under uniform sampling, this results in approximately equal numbers of updates at each setting, matching the per-setting training budget of the fixed-patch suite.
  7. 7.
    Data splits. We use the official train/validation/test split provided by The Well dataset (Ohana et al., 2024). Accordingly, within each dataset, for each set of simulation parameters, we split initial conditions 80/10/10 across train/validation/test. For example, if there are 100 initial conditions for each of 5 parameter settings (each trajectory with 200 timesteps), we use 80 trajectories per parameter setting for training, 10 for validation, and 10 for testing.
C.4 VRMSE metric
The variance scaled mean squared error (VMSE): it is the MSE normalized by the variance of the truth. It has been used for the benchmarking of the Well (Ohana et al., 2024).
We chose to report its square root variant, the VRMSE:
Appendix D Binned spectral error analysis
D.1 Definitions
We assess robustness using the binned spectral normalized mean squared error (BSNMSE) (Ohana et al., 2024). BSNMSE measures the mean squared error between prediction and target after restricting both fields to a spatial frequency band $\mathcal{B}$, and then normalizing by the target energy in that band. Concretely, we first define the (unnormalized) binned spectral mean squared error (BSMSE) as
$\mathrm{BSMSE}_{\mathcal{B}}(u,v)=\left\langle\left|u_{\mathcal{B}}-v_{\mathcal{B}}\right|^{2}\right\rangle,$ (6)
where the band-limited field $u_{\mathcal{B}}$ is obtained by applying a bandpass filter in Fourier space,
[…]
Here, $\mathcal{F}$ denotes the discrete Fourier transform and $\mathbf{1}_{\mathcal{B}}$ is an indicator function selecting frequencies within $\mathcal{B}$.
For each dataset, we define three disjoint frequency bands $\mathcal{B}_{1},\mathcal{B}_{2},$ and $\mathcal{B}_{3}$, corresponding to low, intermediate, and high spatial frequencies. We construct these bands by partitioning wavenumber magnitudes evenly on a logarithmic scale.
Finally, the BSNMSE is defined by normalizing the BSMSE by the target energy in the same band:
$\mathrm{BSNMSE}_{\mathcal{B}}(u,v)=\frac{\left\langle\left|u_{\mathcal{B}}-v_{\mathcal{B}}\right|^{2}\right\rangle}{\left\langle\left|v_{\mathcal{B}}\right|^{2}\right\rangle}.$ (8)
A value of $\mathrm{BSNMSE}_{\mathcal{B}}\geq 1$ indicates that the error at that scale is at least as large as the energy of the target signal, i.e., worse than predicting zero coefficients within band $\mathcal{B}$.
D.2 BSNMSE for different datasets
Fig. 8 shows the next-step prediction BSNMSE of the test set for Rayleigh-Bénard, Turbulent Radiative Layer 2D, Shear Flow  and Supernova Explosion. The spectrum is divided into three frequency bins: low, mid and high being the smallest, intermediate and largest frequency bins whose boundaries are evenly distributed in log space. Crucially, as also emphasized in the main text, we note that the fixed patch models (green) are separately trained models, while CKM (orange) and CSM (blue) (here shown is the flexified vanilla ViT model) are single models trained at randomized strides/patches.
Following our results in Tables 1 and 2 of the main text, our story is reflected in the BSNMSE metric as well. In other words, it is clear from the BSNMSE plot that CSM/CKM (coupled with vanilla ViT for this plot) enable the creation of a single flexible model while maintaining comparable accuracy to the fixed-patch baselines (and often improving it) across these cases. This supports training a single compute-elastic model that can be deployed at multiple inference-time tokenizations.
Figure 8: BSNMSE of the next-step prediction of the test set in the Rayleigh-Bénard, Turbulent Radiative Layer 2D, Shear Flow  and Supernova Explosion datasets. The spectrum is divided into three frequency bins: low, mid and high being the smallest, intermediate and largest frequency bins whose boundaries are evenly distributed in log space. Crucially, as also emphasized in the main text, we note that the fixed patch models (green) are separately trained models, while CKM (orange) and CSM (blue) (here shown is the flexified vanilla ViT model) are single models trained at randomized stride/patch sizes.
Appendix E CViT experiments
CViT experiments.
We further evaluate our flexible patching strategy within the Continuous Vision Transformer (CViT) architecture (Wang et al., 2025b), a recent vision transformer-based model tailored for PDE surrogate modeling. CViT features a patch-based transformer encoder coupled with a novel, grid-based continuous decoder that directly queries spatiotemporal locations. As the decoder operates without patch-based tokenization, we apply our kernel modulation (CKM) only to the encoder, leaving the rest of the architecture unchanged.
A flexified CViT therefore refers to the encoder being a CKM. For CViT (Wang et al., 2025b), whose original design features a single-stage encoder (instead of multi-stage hMLP we used for our vanilla ViT and axial ViT experiments) and a decoder based on grid-based querying, CKM is applied only to the encoder, for the single stage; because the idea is to see if flexification on top of the base fixed patch architecture helps or not. These use cases show CKM’s modularity and adaptability: it can flexify arbitrary convolutional patching pipelines, without requiring architectural changes or any overengineering to the base convolutional or attention mechanism or task head.
Since CViT is designed for 2D problems, we restrict our evaluation to 2D datasets, ensuring a fair and focused test of patch flexibility without altering any other model components. This experiment shows that our flexible tokenization strategy is architecture-agnostic and can be modularly integrated into advanced PDE models like CViT.
Table 8 shows that a flexified CViT consistently outperforms fixed patch size CViT models across four diverse 2D PDE datasets. As in earlier experiments, reducing patch size (e.g., from 16 to 4) increases token count and compute, but yields significantly improved accuracy. Importantly, a flexified CViT provides this accuracy-compute trade-off at inference without retraining, demonstrating the modularity and architectural agnosticism of Overtone. This result also shows how adaptive patching can be extended to complex hybrid architectures like CViT, unlocking new modes of inference-time flexibility in scientific surrogate modeling.
Dataset Tokens/Patch Flexible CViT Fixed CViT
Shear Flow 2048/4 0.120 0.184
128/16 0.127 0.319
Turbulent Radiative Layer 2D 3072/4 0.302 0.299
192/16 0.340 0.364
Active Matter 4096/4 0.054 0.110
256/16 0.065 0.127
In terms of rollouts, since CViT uses a query based continuous decoder, instead of a patch based decoder, it does not inherently suffer from the issues of patch artifacts. Even for that case, Table 9 shows that flexified CViT performs a lot better than the corresponding static patch version. This again shows that our flexification approach is useful even for cases of hybrid encoder-decoder architectures where only the encoder is patch based, but the decoder is not. This reinforces the architecture agnostic-ness and the SOTA performance of our flexible models.